OpenStudy (anonymous):

3(sin2x+cos2x)+3(3−sin2x−cos2x)=28 show that either cos x = 0 or tan x = –1.

OpenStudy (anonymous):

hi! does sin2x mean sin2x or sin^2 x?

OpenStudy (anonymous):

sin2x

OpenStudy (anonymous):

\[3\sin2x + 3\cos2x +3(1-\cos 2x-\sin 2x +2); 3\sin2x +3\cos 2x+3(2\sin^2 x-\sin2x+2); 3\sin2x+3\cos^2x-3\sin^2 x+6\sin^2 x-3\sin2x+6=28; 3\cos^2 x+3\sin^2 x=22; Divide by \cos^2 x, 3+3\tan^2 x= 223\tan^2 x +3= 22+22 \tan^2 x; 19=19\tan^2 x; tan x =-1.\]